OpenStudy (anonymous):

Solve. Round to the nearest hundredth, if necessary. 11/15 = n/180

OpenStudy (anonymous):

nearest hundredth, are you insane???? do you realise how much energy that would take? it would create a black hole and suck us all in! hint, multiply both sides by 180 and then tap what you get into a calculator

OpenStudy (anonymous):

i dont understand

OpenStudy (mayankdevnani):

\[\large \bf \frac{11}{15}=\frac{n}{180}\] Use CROSS MULTIPLICATION RULE,you will get \[\large \bf n=\frac{11 \times 180}{15}\]

OpenStudy (mayankdevnani):

and can you find out the value of n ?? @harz360

OpenStudy (mayankdevnani):